On 1 May, Heloise opened a margin account, buying 400 shares of ABC Company stock for $33 per share, Her initial margin requirement was 55% and the maintenance margin is 25%:

A. At what stock price will Heloise receive a margin call (assuming that Heloise borrowed as much as possible to open an account)?

B. On 1 June, ABC's stock price fell to $15, and Heloise received a margin call. She decides to fulfill the margin requirements by depositing cash in her account. How much cash must Heloise deposit if she is required to restore the margin of 55%.
Answer #1
A) Total value of the purchase = 400*$33 = 13200
Initial margin (borrowing) = 13200*55% = $            7,260
Stock price at which margin call will be received =
P1*400*25% = 7260-(33-P1)*400
where P1 is lowest price for which margin call will
not be made.
Solving for P1
P1*100 = 7260-33*400+P1*400
33*400-7260 = P1*300
P1 = (33*400-7260)/300 = $19.80
Heloise will receive a margin call if the price goes
below $19.80.
B) Margin required at 55% = 400*15*55% = $            3,300
Balance in margin account = 7260-400*(33-15) = $                  60
Cash to be deposited for 55% margin $            3,240